OpenStudy (anonymous):

The base of a rectangular box measures 5 feet by 5 feet. If the volume is 125 cubic feet then the height is how many feet?

OpenStudy (ash2326):

volume= l*b*h l= length b=breadth h=height volume=5*5*h 125=25*h h=5 feet so the rectangular box is a cube.

hero (hero):

V_rect = xyz 125 = 5*5x 125 = 25x 125/25 = x 5 = x